Please show HOW.. preferably with technology..
Answer #1

\\ Find\ X\cdot Y\ and\ X^2

X Y X⋅Y X⋅X
41.9 34.2 1432.98 1755.61
17.8 73.5 1308.3 316.84
8.4 81.2 682.08 70.56
53.4 29.4 1569.96 2851.56
28.5 56.2 1601.7 812.25
-14.6 101.1 -1476.06 213.16
-9.3 74.4 -691.92 86.49
32.6 39.5 1287.7 1062.76
17.3 70.7 1223.11 299.29

\\ Find\ the\ sum\ of\ every\ column\\ \sum{X} = 176 ~,~ \sum{Y} = 560.2 ~,~ \sum{X \cdot Y} = 6937.85 ~,~ \sum{X^2} = 7468.52\\\\ \\ \\ \\ Find\ a\ and\ b\\ \begin{aligned} a &= \frac{\sum{Y} \cdot \sum{X^2} - \sum{X} \cdot \sum{XY} }{n \cdot \sum{X^2} - \left(\sum{X}\right)^2} = \frac{ 560.2 \cdot 7468.52 - 176 \cdot 6937.85}{ 9 \cdot 7468.52 - 176^2} \approx 81.754 \\ \\b &= \frac{ n \cdot \sum{XY} - \sum{X} \cdot \sum{Y}}{n \cdot \sum{X^2} - \left(\sum{X}\right)^2} = \frac{ 9 \cdot 6937.85 - 176 \cdot 560.2 }{ 9 \cdot 7468.52 - \left( 176 \right)^2} \approx -0.998\end{aligned}\\ \\ \\ substitute\ a\ and\ b\ in\ regression \ formula \\ \begin{aligned} y~&=~a ~+~ b \cdot x \\y~&=~81.754 ~-~ 0.998 \cdot x\end{aligned}\\

\\ regression\ equation\ \ \ \boldsymbol{y=81.754-0.998 x}